If Frank and Susan leave their houses at the same time and walk toward
each other, each at his or her own constant rate, they will meet after 5
minutes. If Frank leaves 3 minutes after Susan, they will meet after Frank
has been walking 3 minutes. How long will it take Frank to walk from his
house to Susan’s house?

Answer by Boreal(15235)

In 5 minutes, 5F+5S=1, the whole distance
Also, 3F+6S=1 as well, since Frank walks 3 minutes and Susan 6 minutes to meet.
Set the two equal
5F+5S=3F+6S
S=2F
Therefore, by substitution, 15F=1 or Frank needs 15 minutes to walk there. ANSWER
in 5 minutes, he does 1/3 and Susan 2/3
in 3 minutres, he does 1/5 and Susan 4/5